An electron is to be accelerated in a uniform electric field having a strength of 1.60 x 10 V/m. (a) What energy in keV is given to the electron if it is accelerated through 0.790 m? kev Over what distance would it have to be accelerated to increase its energy by so.o Gev (as in the Stanford Linear Accelerator, which is actually smaller than this)? kon
Answer #1

a) then Energy KE = q(deltaV )

deltaV = E.d = 1.60 * 10^6 * 0.790 = 1264000 V

KE = e * 1264000 V = 1264 keV Answer

b) now KE = 50 GeV = 50 * 10^9 eV

kE = q (Ed)

50 * 10^9 eV = 1e * 1.60 * 10^6 V/m * L

L = 31250 m = 31.25 km Answer

I hope help you !!